# RLHF Workflow: From Reward Modeling to Online RLHF

### v1

@@ Abstract
We present the workflow of Online Iterative Reinforcement Learning from Human Feedback (RLHF) in this technical report, which is widely reported to outperform its offline counterpart by a large margin in the recent large language model (LLM) literature. However, existing open-source RLHF projects are still largely confined to the offline learning setting. In this technical report, we aim to fill in this gap and provide a detailed recipe that is easy to reproduce for online iterative RLHF. In particular, since online human feedback is usually infeasible for open-source communities with limited resources, we start by constructing preference models using a diverse set of open-source datasets and use the constructed proxy preference model to approximate human feedback. Then, we discuss the theoretical insights and algorithmic principles behind online iterative RLHF, followed by a detailed practical implementation. Our trained LLM, SFR-Iterative-DPO-LLaMA-3-8B-R, achieves impressive performance on LLM chatbot benchmarks, including AlpacaEval-2, Arena-Hard, and MT-Bench, as well as other academic benchmarks such as HumanEval and TruthfulQA. We have shown that supervised fine-tuning (SFT) and iterative RLHF can obtain state-of-the-art performance with fully open-source datasets. Further, we have made our models, curated datasets, and comprehensive step-by-step code guidebooks publicly available. Please refer to https://github.com/RLHFlow/RLHF-Reward-Modeling and https://github.com/RLHFlow/Online-RLHF for more detailed information.
@@ 2.1 Preference Datasets
  • Distilabel-Orca [^3] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs is collected similarly with Capybara but with the prompts from Lian et al., (2023).
@@ 2.3 Evaluation Result
. Base Model Type Data Mixture Chat Chat Hard Safety Reasoning LLaMA-3-8B-it Prompting - 93.6 44.3 71.3 73.5 LLaMA-2-13B BT mix1 96.4 55.5 55.0 62.4 LLaMA-3-8B-it BT mix2 99.4 65.0 87.7 87.8 LLaMA-3-8B-it Preference mix2 98.6 65.8 89.6 94.9 Gemma-2B-it Prompting - 70.3 42.3 38.2 50.0 Gemma-2B-it BT mix2 95.0 40.8 81.2 74.2 Gemma-2B-it Preference mix2 96.0 40.5 82.8 80.7
@@ 3.1 Supervised Fine-Tuning
The base model used in this project is LLaMA-3-8B. To ensure the reproducibility and openness of the project, we perform SFT by ourselves to obtain the initial policy $\pi_{0}$. We collect a set of high-quality instruction datasets for SFT, such as ShareGPT, SlimOrca, MathInstruct, and Evol-Instruct (see the Appendix for a full list). The training is carried out for one epoch with a learning rate of $2\times 10^{-5}$. A cosine scheduler is employed, and the global batch size is set to 32 with a warm-up ratio of 0.03. To accelerate training, we follow Diao et al., (2023); Tunstall et al., (2023) to pack the samples and use a block size of 8192. We use the SFT-Trainer from TRL and the detailed recipe is available at https://github.com/RLHFlow/Online-RLHF.
@@ 3.3 Practical Implementation Details
Prompt set, and data generation. We collect prompts from UltraFeedback (Cui et al.,, 2023), HelpSteer (Wang et al.,, 2023), OpenOrca (Lian et al.,, 2023), UltraInteract (Yuan et al., 2024a, ), Capybara (Daniele and Suphavadeeprasit,, 2023) and DIBT-10K [^6] https://huggingface.co/datasets/DIBT/10k_prompts_ranked and prepare the full prompt set here [^7] https://huggingface.co/datasets/RLHFlow/prompt-collection-v0.1. In our experiments, we use a subset of 60K prompts and iterate for three iterations, so 20K prompts are used to generate 20K x 8 responses per iteration. To accelerate data generation, we use VLLM (Kwon et al.,, 2023) for inference. We set the max generation length as 2048, and use a sampling temperature of 1.0/0.7 without any top-k/top-p strategy. To offer a more intuitive comprehension of our prompts collection, we provide visualization plots in Figure 7, which are generated on Nomic Atlas [^8] https://atlas.nomic.ai/ with the nomic-embed-text-v1.5 text embedding model (Nussbaum et al.,, 2024).
@@ 4.2 Main Results
Ours (Iterative RLHF) 8B Iterative DPO 37.2 8.46 29.1
@@
Online iterative RLHF significantly improves conversation quality. We evaluate our model’s conversation abilities using AlpacaEval-2, MT-Bench, and Chat-Arena-Hard, (results in Table 4). Compared to other open-source models with less than 10B parameters, our model – SFR-Iterative-DPO-LLaMA-3-8B-R outperforms them on the conversation and instruction-following benchmarks with a significant margin. Notably, our model trained with iterative DPO consistently outperforms that of vanilla offline DPO (DPO baseline). This demonstrates the advantage of online iterative RLHF. Moreover, our model outperforms the Tulu-2-DPO-70B and GPT-3.5-turbo-1106, which are aligned by DPO or PPO and are much larger than our base model. These results show that the online iterative RLHF can effectively adjust the style of the model responses, thus improving the conversation quality.
@@
The RLHF-aligned models based on some initial checkpoints and more epochs can approach or even outperform the state-of-the-art closed-source models like GPT-4 and Claude on benchmarks. However, we remark that we should be more careful in interpreting these results because the test sets of the benchmarks are finite and may not be representative enough to capture the complicated real-world scenarios. Moreover, the increased possibility of small models overfitting the benchmarks may lead to benchmark hacking, which means that the real capacity of a model with a high score is still limited. In particular, while it is possible to get even higher results on the benchmark (e.g., 44.84 in LC AlpacaEval-2 and 35.7 in Chat-Arena-hard, but the performance on academic benchmarks drops significantly), we presented our current model by human evaluation on some randomly chosen test prompts.
@@
Ours - 37.2 29.1 656 80.7 65.3 64.6 62.2 64.3 60.8
Ours-concise 0.001 41.9 22.1 382 78.8 65.5 66.5 60.4 65.1 62.4
@@ 5 End Note and Future Direction
In this technical report, we study the workflow of the online iterative RLHF, which leverages on-policy sampling and external preference signals from a proxy preference model trained on a diverse set of open-source preference datasets. The resulting model (referred to as SFR-Iterative-DPO-LLaMA-3-8B-R) demonstrates impressive performance on standard benchmarks, and the report provides detailed instructions for reproducing the results, including data, code, models, and hyper-parameter choices.
@@
6 Acknowledgements
The authors would like to thank the great open-source communities, including the Huggingface TRL team, the Huggingface H4 team, the Allen Institute AI RewardBench team, the Meta LLaMA team, and Axolotl team for sharing the models, codes, and training sets. We also thank Zihao Li for assistance in reward modeling.

### v2

@@ Abstract
We present the workflow of Online Iterative Reinforcement Learning from Human Feedback (RLHF) in this technical report, which is widely reported to outperform its offline counterpart by a large margin in the recent large language model (LLM) literature. However, existing open-source RLHF projects are still largely confined to the offline learning setting. In this technical report, we aim to fill in this gap and provide a detailed recipe that is easy to reproduce for online iterative RLHF. In particular, since online human feedback is usually infeasible for open-source communities with limited resources, we start by constructing preference models using a diverse set of open-source datasets and use the constructed proxy preference model to approximate human feedback. Then, we discuss the theoretical insights and algorithmic principles behind online iterative RLHF, followed by a detailed practical implementation. Our trained LLM, LLaMA-3-8B-SFR-Iterative-DPO-R, achieves impressive performance on LLM chatbot benchmarks, including AlpacaEval-2, Arena-Hard, and MT-Bench, as well as other academic benchmarks such as HumanEval and TruthfulQA. We have shown that supervised fine-tuning (SFT) and iterative RLHF can obtain state-of-the-art performance with fully open-source datasets. Further, we have made our models, curated datasets, and comprehensive step-by-step code guidebooks publicly available. Please refer to https://github.com/RLHFlow/RLHF-Reward-Modeling and https://github.com/RLHFlow/Online-RLHF for more detailed information.
@@ 2.1 Preference Datasets
  • Distilabel-Orca [^3] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs is collected similarly with Capybara but with the prompts from Lian et al., 2023a .
@@ 2.3 Evaluation Result
. Base Model Type Data Mixture Chat Chat Hard Safety Reasoning LLaMA-3-8B-it Prompting - 93.6 44.3 71.3 73.5 LLaMA-2-13B BT mix1 96.4 55.5 55.0 62.4 LLaMA-3-8B-it BT mix2 99.4 65.1 87.8 86.4 LLaMA-3-8B-it Preference mix2 98.3 65.8 89.7 94.7
@@ 3.1 Supervised Fine-Tuning
The base model used in this project is LLaMA-3-8B. To ensure the reproducibility and openness of the project, we perform SFT by ourselves to obtain the initial policy $\pi_{0}$. We collect a set of high-quality instruction datasets for SFT, such as ShareGPT, SlimOrca (Lian et al., 2023b, ), MathInstruct (Yue et al.,, 2023), and Evol-Instruct (see the Appendix for a full list). The training is carried out for one epoch with a learning rate of $2\times 10^{-5}$. A cosine scheduler is employed, and the global batch size is set to 32 with a warm-up ratio of 0.03. To accelerate training, we follow Diao et al., (2023); Tunstall et al., (2023) to pack the samples and use a block size of 8192. We use the SFT-Trainer from TRL and the detailed recipe is available at https://github.com/RLHFlow/Online-RLHF.
@@ 3.3 Practical Implementation Details
Prompt set, and data generation. We collect prompts from UltraFeedback (Cui et al.,, 2023), HelpSteer (Wang et al.,, 2023), OpenOrca (Lian et al., 2023a, ), UltraInteract (Yuan et al., 2024a, ), Capybara (Daniele and Suphavadeeprasit,, 2023) and DIBT-10K [^6] https://huggingface.co/datasets/DIBT/10k_prompts_ranked and prepare the full prompt set here [^7] https://huggingface.co/datasets/RLHFlow/prompt-collection-v0.1. In our experiments, we use a subset of 60K prompts and iterate for three iterations, so 20K prompts are used to generate 20K x 8 responses per iteration. To accelerate data generation, we use VLLM (Kwon et al.,, 2023) for inference. We set the max generation length as 2048, and use a sampling temperature of 1.0/0.7 without any top-k/top-p strategy. To offer a more intuitive comprehension of our prompts collection, we provide visualization plots in Figure 7, which are generated on Nomic Atlas [^8] https://atlas.nomic.ai/ with the nomic-embed-text-v1.5 text embedding model (Nussbaum et al.,, 2024).
@@ 4.2 Main Results
Ours (Iterative RLHF) 8B Iterative DPO 31.3 8.46 29.1
@@
Online iterative RLHF significantly improves conversation quality. We evaluate our model’s conversation abilities using AlpacaEval-2, MT-Bench, and Chat-Arena-Hard, (results in Table 4). Compared to other open-source models with less than 10B parameters, our model – LLaMA-3-8B-SFR-Iterative-DPO-R outperforms them on the conversation and instruction-following benchmarks with a significant margin. Notably, our model trained with iterative DPO consistently outperforms that of vanilla offline DPO (DPO baseline). This demonstrates the advantage of online iterative RLHF. Moreover, our model outperforms the Tulu-2-DPO-70B and GPT-3.5-turbo-1106, which are aligned by DPO or PPO and are much larger than our base model. These results show that the online iterative RLHF can effectively adjust the style of the model responses, thus improving the conversation quality.
@@
The RLHF-aligned models based on some initial checkpoints and more epochs can approach or even outperform the state-of-the-art closed-source models like GPT-4 and Claude on benchmarks. However, we remark that we should be more careful in interpreting these results because the test sets of the benchmarks are finite and may not be representative enough to capture the complicated real-world scenarios. Moreover, the increased possibility of small models overfitting the benchmarks may lead to benchmark hacking, which means that the real capacity of a model with a high score is still limited. In particular, while it is possible to get even higher results on the benchmark (e.g., 44.84 in LC AlpacaEval-2 and 35.7 in Chat-Arena-hard, but the performance on academic benchmarks drops significantly), we presented our current model by human evaluation on some randomly chosen test prompts. We also found that GPT-based evaluation highly depends on the configuration. Our model obtains 37.2 LC win-rate (45.4 win-rate) with “alpaca_eval_gpt4_turbo_fn” config, which has better agreement with human evaluation.
@@
Ours - 31.2 29.1 656 80.7 65.3 64.6 62.2 64.3 60.8
Ours-concise 0.001 38.1 22.1 382 78.8 65.5 66.5 60.4 65.1 62.4
@@ 5 End Note and Future Direction
In this technical report, we study the workflow of the online iterative RLHF, which leverages on-policy sampling and external preference signals from a proxy preference model trained on a diverse set of open-source preference datasets. The resulting model (referred to as LLaMA-3-8B-SFR-Iterative-DPO-R) demonstrates impressive performance on standard benchmarks, and the report provides detailed instructions for reproducing the results, including data, code, models, and hyper-parameter choices.
@@ Appendix B Additional Experimental Details
Evaluation. For MT-Bench evaluation, we use huggingface default generation and “You are CMB, an AI assistant known for its intelligence and expertise across all fields of knowledge. You are designed to provide detailed and helpful responses to a wide range of user inquiries, ensuring clarity and accuracy in every interaction.” as system prompt. Transformers version is 4.38.1. For AlpacaEval and Chat-Arena-Hard, we use vllm to generate samples.